Gas expands at constant pressure p=6,2.10⁵ Pa from volume V₁=3,2.10⁻³ m³ to volume V₂=5,2,10⁻³m³. How much work has been done by the gas?


1
Expert's answer
2021-03-02T18:04:26-0500

By the definition of the work done by gas, we have:


"W=p\Delta V=p(V_2-V_1),""W=6.2\cdot10^5\ Pa\cdot(5.2\cdot10^{-3}\ m^3-3.2\cdot10^{-3}\ m^3)=1240\ J."